# Lumos : Empowering Multimodal LLMs with Scene Text Recognition

### v1

@@ 1. Introduction
Now assuming we choose an on-device STR solution, the second challenge is the constrained compute and memory resources on devices. Although running STR models on-device may seem like a viable solution to address latency concerns, current state-of-the-art STR models are not readily suitable for on-device usage; for example, Google’s recent work (Long et al., 2022) features a text detection model that alone has a size of 240MB, impractical for on-device use where several other processes might be running and sharing memory.
@@ 4.2. Text Detection
Problem definition and challenges Text detection takes the cropped image in full-resolution as input, predicts location of each word as bounding boxes. There are three challenges for detecting text in the wild: C1. the text size can be very small (e.g., ingredients on a coke can at arm length) or very big (e.g., storefront); C2. text can often be tilted with the nature of the image; C3. we are not able to use state-of-the-art text detection model architectures like (Liao et al., 2019; Long et al., 2022) with the on-device constraint.
@@ 6.1. Experiment Setup
Datasets Table 1 summarizes the datasets we used for evaluation. We have two benchmarks: In-house wild text benchmark and Public wild text benchmark. In-house wild text benchmark contains 968 in-the-wild text images taken from an edge device and contains 47K word boxes. The benchmark contains annotations for the word boxes and transcriptions, and in addition annotations for salient areas for ROI evaluation. Public wild text benchmark is a broadly-used STR benchmark, containing 1.7K images and 146K word boxes.
@@
To understand the quality of STR solutions, we measured the Word Error Rate (WER), a standard metric extensively used in the domain of speech and text recognition. WER considers 3 types of errors: 1) Deletion: a ground truth word that is not detected; 2) Insertion: a prediction that is not matched to any ground truth word box; 3) Substitution: a prediction that matches a ground truth box, but the word recognized is different from the ground truth. WER is the sum of Deletion, Insertion, Substitution errors divided by the total number of words in the ground truth. With the existence of insertion errors, WER can be higher than 1. A lower WER is indicative of higher quality of the models.
@@ 6.2. End-to-End Quality
Public wild text benchmark
@@
Google OCR (goo, [n. d.]) 30.4% 9.4% 9.5% 11.5% 240Mb+ [^2] Estimated based on (Long et al., 2022), using the size of MaX-DeepLab-S (Wang et al., 2021)

### v2

@@ 1. Introduction
Now assuming we choose an on-device STR solution, the second challenge is the constrained compute and memory resources on devices. Although running STR models on-device may seem like a viable solution to address latency concerns, current state-of-the-art STR models are not readily suitable for on-device usage; for example, Google’s recent work (Long et al., 2022a) features a text detection model that alone has a size of 240MB, impractical for on-device use where several other processes might be running and sharing memory.
@@ 4.2. Text Detection
Problem definition and challenges Text detection takes the cropped image in full-resolution as input, predicts location of each word as bounding boxes. There are three challenges for detecting text in the wild: C1. the text size can be very small (e.g., ingredients on a coke can at arm length) or very big (e.g., storefront); C2. text can often be tilted with the nature of the image; C3. we are not able to use state-of-the-art text detection model architectures like (Liao et al., 2019; Long et al., 2022a) with the on-device constraint.
@@ 6.1. Experiment Setup
Datasets Table 1 summarizes the datasets we used for evaluation. We have two benchmarks: In-house wild text benchmark and Public wild text benchmark (Long et al., 2022b). In-house wild text benchmark contains 968 in-the-wild text images taken from an edge device and contains 47K word boxes. The benchmark contains annotations for the word boxes and transcriptions, and in addition annotations for salient areas for ROI evaluation. Public wild text benchmark is a broadly-used STR benchmark, containing 1.7K images and 146K word boxes.
@@
To understand the quality of STR solutions, we measured the Word Error Rate (WER), a standard metric extensively used in the domain of speech (Shenoy et al., 2021b, a; Bekal et al., 2021) and text recognition (Dingliwal et al., 2022, 2021). WER considers 3 types of errors: 1) Deletion: a ground truth word that is not detected; 2) Insertion: a prediction that is not matched to any ground truth word box; 3) Substitution: a prediction that matches a ground truth box, but the word recognized is different from the ground truth. WER is the sum of Deletion, Insertion, Substitution errors divided by the total number of words in the ground truth. With the existence of insertion errors, WER can be higher than 1. A lower WER is indicative of higher quality of the models.
@@ 6.2. End-to-End Quality
Public wild text benchmark (Long et al., 2022b)
@@
Google OCR (goo, [n. d.]) 30.4% 9.4% 9.5% 11.5% 240Mb+ [^2] Estimated based on (Long et al., 2022a), using the size of MaX-DeepLab-S (Wang et al., 2021)